# Unified Visual Understanding and Generation

### v3

@@ 4.1 Main Results
MagicBrush [56] × × × × × × 1.85 2.72 1.47 1.31 1.89 1.57 1.80
Instruct-P2P [3] × × × × × × 1.89 2.29 1.79 1.33 1.93 1.49 1.48
AnyEdit [49] × × × × × × 2.63 3.12 2.66 1.82 2.71 2.34 2.07
Step1-Edit [27] × × × × × × 3.17 3.90 3.13 1.87 3.45 2.61 2.52
@@ 4.3 Image Manipulation
As illustrated in Table 4, we compare UniWorld-V1 with other open-source models and GPT-4o-Image on editing capabilities, using the ImgEdit-Bench [48] benchmark. Our UniWorld-V1 model demonstrates the best overall performance among all open-source models, achieving an impressive total score of 3.37. This significantly surpasses other leading open-source models such as Step1X-Edit (3.17) and BAGEL (3.17), fully showcasing UniWorld-V1’s exceptional and versatile performance across diverse image editing domains. Notably, UniWorld-V1 achieves the highest scores among all open-source models in several key categories: Adjust (3.70), Remove (3.54), Extract (2.23), Replace (3.49), and Hybrid (3.13). These leading scores underscore UniWorld-V1’s advanced capabilities in precise attribute adjustment, specified element removal, object extraction, hybrid editing, and content replacement. While GPT-4o-Image maintains a leading position with a total score of 4.31, UniWorld-V1’s total score of 3.37 makes it the closest performing model to GPT-4o-Image among all open-source alternatives. This strongly demonstrates UniWorld-V1’s significant step towards achieving image editing capabilities comparable to industry-leading models.
@@
[…]
MagicBrush [56] 2.72 1.47 1.31 1.89 1.57 2.49 1.39 1.80 2.03 1.85
Instruct-P2P [3] 2.29 1.79 1.33 1.93 1.49 3.54 1.51 1.48 1.67 1.89
AnyEdit [49] 3.12 2.66 1.82 2.71 2.34 3.27 3.31 2.07 2.37 2.63
UltraEdit [59] 3.63 3.01 2.02 3.13 1.71 3.69 3.57 2.33 3.31 2.93
Step1X-Edit [27] 3.90 3.13 1.87 3.45 2.61 4.44 3.43 2.52 3.19 3.17
BAGEL [9] 3.55 3.30 1.56 3.38 2.44 4.24 4.29 2.55 3.22 3.17
GPT-4o-Image 4.65 4.26 2.96 4.49 3.81 4.75 4.76 4.54 4.62 4.31
UniWorld-V1 3.86 3.70 2.23 3.49 3.54 4.22 3.44 3.13 2.76 3.37

### v4

@@ 4.1 Main Results
MagicBrush [56] × × × × × × 1.83 2.84 1.58 1.51 1.97 1.58 1.62
Instruct-P2P [3] × × × × × × 1.88 2.45 1.83 1.44 2.01 1.50 1.20
AnyEdit [49] × × × × × × 2.45 3.18 2.95 1.88 2.47 2.23 1.56
UltraEdit [59] × × × × × × 2.70 3.44 2.81 2.13 2.96 1.45 1.91
Step1X-Edit [27] × × × × × × 3.06 3.88 3.14 1.76 3.40 2.41 2.64
@@ 4.3 Image Manipulation
As illustrated in Table 4, we compare UniWorld-V1 with other open-source models and GPT-4o-Image on editing capabilities using the ImgEdit-Bench [48] benchmark. Our UniWorld-V1 model achieves the best overall performance among all open-source models, with a total score of 3.26, outperforming other strong open-source competitors such as BAGEL (3.20) and Step1X-Edit (3.06). This highlights UniWorld-V1’s robust and consistent image editing ability across a wide spectrum of tasks. Notably, UniWorld-V1 ranks first among open-source models in several key categories, including Adjust (3.64), Extract (2.27), Replace (3.47), Remove (3.24), Background (2.99), and Hybrid (2.96). These results reflect UniWorld-V1’s strong capability in tasks that require nuanced attribute adjustment, structural modification, and complex multi-operation editing. While GPT-4o-Image continues to lead with a remarkable total score of 4.20, UniWorld-V1 emerges as the closest open-source contender, significantly narrowing the performance gap. This demonstrates UniWorld-V1’s notable progress towards achieving high-fidelity and generalizable image editing quality, comparable to state-of-the-art proprietary models.
Table 4: Comparison results on ImgEdit-Bench. “Overall” is calculated by averaging all scores across tasks. We use GPT-4.1 for evaluation.
@@
MagicBrush [56] 2.84 1.58 1.51 1.97 1.58 1.75 2.38 1.62 1.22 1.83
Instruct-P2P [3] 2.45 1.83 1.44 2.01 1.50 1.44 3.55 1.20 1.46 1.88
AnyEdit [49] 3.18 2.95 1.88 2.47 2.23 2.24 2.85 1.56 2.65 2.45
UltraEdit [59] 3.44 2.81 2.13 2.96 1.45 2.83 3.76 1.91 2.98 2.70
Step1X-Edit [27] 3.88 3.14 1.76 3.40 2.41 3.16 4.63 2.64 2.52 3.06
BAGEL [9] 3.56 3.31 1.70 3.30 2.62 3.24 4.49 2.38 4.17 3.20
GPT-4o-Image 4.61 4.33 2.90 4.35 3.66 4.57 4.93 3.96 4.89 4.20
UniWorld-V1 3.82 3.64 2.27 3.47 3.24 2.99 4.21 2.96 2.74 3.26